# Implementation of an ITER-relevant QP-based Current Limit Avoidance algorithm in the TCV tokamak

### v1

@@ 4 Experimental results
The effectiveness of the proposed CLA approach is demonstrated in three different test cases on a TCV lower single-null reference scenario, shown in Fig. 2:
  • Test case I: the limit on one circuit is fictitiously reduced.
  • Test case II: the limits on three circuits are fictitiously reduced.
  • Test case III: the limit on one circuit is fictitiously reduced in a time-varying fashion.
@@
Figure 4: Reference scenario - time traces of the isoflux control errors. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@
The real-time implementation of the CLA system relies on the availability of an effective solver for the QP problem (4). The implementation proposed for this work is based on the C++ library [28], which implements the Goldfarb-Idnani active-set dual method [29]. As it can be seen from figs.10-15-20, the solver converged within 3 iterations for all the considered cases, even when multiple current saturations were hit simultaneously.
@@ 4.1 Test Case I
Figure 7: Test Case I - time traces of the isoflux control errors. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@ 4.2 Test Case II
Figure 12: Test Case II - time traces of the isoflux control errors. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@ 4.3 Test Case III
The results for the third test case, TCV shot #83438, are reported in figs. 16-20. For this case, the saturation on the E5 coils decreases in steps of $250$ A applied every $0.2$ s. The snapshots at three different time instants following the CLA activation are shown in Fig. 16, while the time traces of the shape control errors are shown in Fig. 18. The shape accuracy degradation is comparable to the one observed in test case II, but in this case the time-varying nature of the saturation limit is clearly reflected by the small jumps observed in the shape control error traces in Fig. 18 and in the behavior of the QP cost function, reported in Fig. 20. The relevant coil currents and corresponding saturation limits are reported in Fig. 18, where the variable saturation limit for the E5 current is also shown. The E5 current returns into the acceptable time-varying band following the activation of the CLA at $t=0.6$. The iterations performed by the QP solver are shown in Fig. 20.
@@
Figure 17: Test Case III - time traces of the isoflux control errors. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@ 5 Conclusive Remarks
This paper describes the implementation of an effective solution to tackle the problem of the PF currents saturation. The methodology proposed for inclusion in the ITER PCS has been tested experimentally on the TCV tokamak. The algorithm proved to be capable of dealing robustly with the saturation of several active circuit currents, up to four in Test Case II, and also proved to be capable of handling time-varying saturations, as demonstrated in Test Case III. The simplicity of the proposed scheme and the tractable computation cost make this algorithm an effective tool for future reactors such as ITER.

### v2

@@ 4 Experimental results
The effectiveness of the proposed CLA approach is demonstrated in three different test cases on a TCV lower single-null reference scenario, shown in Fig. 2. For all the considered cases, saturations of ±$\pm 3.5$ kA have been imposed on all E and F coils. Moreover, the following situations have been considered:
  • Test case I: the limit on one circuit is fictitiously reduced to $2.5$ kA.
  • Test case II: the limits on three circuits are fictitiously reduced to $2.5$ kA.
  • Test case III: like test case II, but the limit on one of the circuits is fictitiously reduced in a time-varying fashion.
@@
Figure 4: Reference scenario (TCV pulse #79742) - time traces of the isoflux control errors. The poloidal flux errors (at contour and strike points) are shown in blue, while the magnetic field errors (at the X-point) are in red. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@
With reference to the optimization problem (3), diagonal $W$ and $Q$ weighing matrices have been used. The individual weights in $W$ can be used to promote accuracy on specific channels, for instance on the control of radial gaps for configurations which are very close to the inner or outer walls. In the proposed examples, the same weights used in the shape controller were kept, i.e. a weight of $2$ on the strike points and of $1$ on all the other channels. The weight for each control point is then scaled proportionally to the inverse of the gradient of the associated controlled quantity, to “equalize” the shape accuracy in terms of local boundary deformations.
For what concerns the weights in $Q$ instead, a unitary gain was simply chosen on all channels. A further scalar relative weight between $W$ and $Q$ can be introduced to empirically regulate the time scale on which the other currents are rearranged to compensate for the ones affected by the hard constraints. In practice, a lower weight on $Q$ relative to $W$ means that the allocator has more freedom to modify the requested currents in order to minimize the projected shape error. A first guess for such a scaling can be obtained by looking at the values of the typical shape control errors and of the reference currents in a similar discharge, but then it is usually tuned empirically based on offline simulations. For the proposed examples, we multiplied the original $W$ matrix by a factor ×$\times 10^{3}$.
The real-time implementation of the CLA system relies on the availability of an effective solver for the QP problem (3). The implementation proposed for this work is based on the C++ library [30], which implements the Goldfarb-Idnani active-set dual method [31]. As it can be seen from figs.10-15-22, the solver converged within 3 iterations for all the considered cases, even when multiple current saturations were hit simultaneously.
@@ 4.1 Test Case I
Figure 7: Test Case I (TCV pulse #83435) - time traces of the isoflux control errors. The poloidal flux errors (at contour and strike points) are shown in blue, while the magnetic field errors (at the X-point) are in red. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@ 4.2 Test Case II
Figure 12: Test Case II (TCV pulse #83437) - time traces of the isoflux control errors. The poloidal flux errors (at contour and strike points) are shown in blue, while the magnetic field errors (at the X-point) are in red. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@
We take advantage of this example to illustrate how the steady-state accuracy of the underlying control loop affects the precision of the imposed saturation limits, as discussed in section 3. Consider the currents in coils E005 and E006. These coils are not used for plasma position control. Hence, their currents are directly controlled by hybrid and the respective entries in the static-gain matrix $P_{EF}^{0}$ are approximately unitary from reference to corresponding current and zero elsewhere, i.e. for these control channels $P^{0}_{EF}r_{h}\approx r_{h}$.
Fig. 16 shows the currents in these coils for both the “baseline” scenario (pulse #79742, dashed black) and the considered test case (green). It can be seen how the current on E005 is significantly reduced with respect to the reference case, and successfully brought below the saturation threshold. The current in E006 is also reduced, but it is still slightly above the saturation limit (by approximately $50$ A, i.e. $\sim 2\%$ of its value). However, the projected steady-state reference value $P_{EF}^{0}(r_{h}+u_{sh}+u_{CLA})$ is correctly lower than the imposed saturation limit of $2500$ A. The same is true for the $I_{EF}^{0}$ quantity, which is approximately equal to the saturation limit (minus the effect of high frequency noise, coming from the measured $I_{EF}$ values in (1). Similarly, for coil E005, the expected $I_{EF}^{0}$ value (blue trace) is exactly at the saturation limit, while the achieved current has a small offset, this time in a beneficial direction, of approximately $150$ A ($\sim 6\%$). As discussed above, for both cases the effect of the projection matrix $P_{EF}^{0}$ is negligible. As a consequence, the residual error between the modified reference $r_{h}+u_{sh}+u_{CLA}$ (red trace) and the achieved current (green trace) should be attributed to the accuracy of the hybrid controller.
For completeness, in Fig. 17 the effects of the different terms in (1) are considered. In particular, the panels show the (projected) E005 and E006 current errors with the allocator and shape controller actions (dotted black lines). Moreover, the individual contributions to the modified references due to the shape control and the allocator are shown in red and green, respectively. Again, it is worth to remark that the action of the effect of the projector $P_{EF}^{0}$ is negligible for these currents, hence these traces can be considered in practice as error or reference signals on the coil currents themselves. Looking at the case of E005 (left plot), it can be seen for instance how the allocator introduces an offset at steady-state of about $-950$ A, which brings the current back into the saturation limits. Moreover, the error between the modified reference and the hybrid feedback signal is reported in black. Its nonzero steady-state value is also reflected in the small offset between the green and red traces in Fig. 16
Figure 16: TCV pulse #83437. In green the measured $I_{EF}$ currents, in blue the projected steady-state values $I_{EF}^{0}$, in red the reference to the hybrid controller modified by the action of the shape control and of the CLA system and projected through the steady-state gain $P_{EF}^{0}$. In black the same currents in the reference pulse #79742.
Figure 17: TCV pulse #83437. In black the tracking error on the currents in the E005 (left) and E006 (right) coils. In green and red the reference modifications introduced by the shape controller and the allocator respectively.
@@ 4.3 Test Case III
The results for the third test case, TCV shot #83438, are reported in figs. 18-22. For this case, the saturation on the E5 coils decreases in steps of $250$ A applied every $0.2$ s. The snapshots at three different time instants following the CLA activation are shown in Fig. 18, while the time traces of the shape control errors are shown in Fig. 20. The shape accuracy degradation is comparable to the one observed in test case II, but in this case the time-varying nature of the saturation limit is clearly reflected by the small jumps observed in the shape control error traces in Fig. 20 and in the behavior of the QP cost function, reported in Fig. 22. The relevant coil currents and corresponding saturation limits are reported in Fig. 20, where the variable saturation limit for the E5 current is also shown. The E5 current returns into the acceptable time-varying band following the activation of the CLA at $t=0.6$. The iterations performed by the QP solver are shown in Fig. 22.
It should be noted that in general time-varying saturations are not expected in tokamaks; in fact, this example was mainly designed as a “stress-test” to showcase the flexibility of the proposed architecture. However, a similar case of practical interest may occur if one wished to introduce constraints on currents that are already beyond the saturation level at the activation time of the CLA system in a gradual fashion, to avoid large transients due to these currents being brought abruptly back into the saturation limits. It can be noticed how, for the considered case, the peak transient shape error in Fig. 20 is indeed reduced compared to Figs. 8 and 13.
@@
Figure 19: Test Case III (TCV pulse #83438) - time traces of the isoflux control errors. The poloidal flux errors (at contour and strike points) are shown in blue, while the magnetic field errors (at the X-point) are in red. The beginning and end of the controlled time window are indicated by the dashed black lines.
@@ 5 Conclusive Remarks
This paper describes the implementation of an effective solution to tackle the problem of the PF coil currents saturation. The methodology proposed for inclusion in the ITER PCS has been tested experimentally on the TCV tokamak. The algorithm proved to be capable of robustly dealing with the saturation of several active circuit currents, as shown by the example of section 4.2, and also proved to be capable of handling time-varying saturations, as demonstrated by the example discussed in section 4.3.
It is worth to remark that since the proposed CLA strategy is only related to the expected steady-state behavior, it does not provide any performance guarantee during transients. In fact, the introduction of hard constraints on coil currents could lead to fast variations in the latter ones, which could potentially perturb the plasma configuration. One possibility to mitigate this could be that of introducing the saturation limits gradually, for example as demonstrated by the third proposed example.
However, it should also be noted that the transient performance depends fundamentally on the underlying control architecture, and is not affected by the introduction of the allocator, which only modifies the reference signals to such controller. As it can be seen from the results in section 4, for the considered cases the observed shape degradation during transients remained on an acceptable level.
As discussed in the introduction to this article, a control technique that is often compared to current allocation algorithms in terms of its capability to handle hard constraints is MPC. Compared to the proposed CLA, MPC has the advantage of offering an integrated solution to deal with both control requirements and constraints, and it naturally provides finer control over transient phases. However, it should be noted that even with MPC, if a hard constraint is activated on a coil whose current is already significantly beyond the saturation threshold, a fast transient is inevitable.
The proposed CLA, on the other hand, offers two main advantages over MPC. First, since it only deals with steady-state constraints, its implementation is much lighter from a computational point of view. Moreover, especially when a coil current controller is installed on the tokamak, a detailed knowledge of the underlying control architecture or of the plasma dynamics is not necessary.
The simplicity of the proposed scheme and the tractable computation cost make this algorithm an effective tool for future reactors such as ITER.